# MACEval: A Multi-Agent Continual Evaluation Network for Large Models

### v1

@@ Appendix E Future Possibilities
Apart from the comments for possible next steps of research related to the evaluation of large models that have already been given, this section is devoted to the extension for some of them and then more topics with good potential based upon our understanding and rethinking of the field.
@@
  • Evaluating the evaluations. Evaluations are critical for understanding the capabilities of large models. Its fairness, comprehensiveness, reproducibility, timeliness, scalability, transparency, and practicality are key points to serve as a "good" evaluation. Recently, with the rapidly increasing number of various benchmarks, it is highly needed for objective metrics to evaluate evaluations themselves.
@@
Appendix F Prompt Templates
F.1 Prompt Templates for Visual Perception Tasks
@@
<|System Prompt|> You are a helpful assistant for prompt generation in text-to-image tasks.
<|User Prompt|> Please generate a prompt regarding the generation of {keyword} images. Return text in <text></text> format.
<|Interviewer’s Answer|> <text>A high-resolution natural landscape at sunrise, viewed from a hilltop. The foreground shows dewy grass and wildflowers gently swaying in the breeze. In the middle ground, a winding dirt path leads through a lush green forest with scattered rays of sunlight filtering through the trees. In the background, soft morning mist blankets a valley with a river glistening under the golden light of the rising sun. Birds are flying low near the horizon, and the sky transitions from deep orange to pale blue.</text>
<|External image generation model|> def SD3_5 (prompt):
<|External distortion algorithm|> def add_noise (img, noise_level):
<|Resulting content|> Reference image: <image1>
@@
<|User Prompt|> Given two images <image1> and <image2>, please answer if there are any differences in noise between them. Return judgment <yes> or <no> with analysis in format:
@@
<|Interviewee’s Answer|> <Judgment>No</Judgment>
@@
F.2 Prompt Templates for Content Understanding Tasks
@@
<|System Prompt|> You are a helpful assistant.
<|User Prompt|> Please generate two random numbers from 1 to 100 and from 1 to {{level+2} $\times$ {level+2}}, respectively. Return answer in the following format:
@@
<|Interviewer’s Answer|> Number1: <67>
@@
<|External grid_image generation algorithm|> def grid_image (image_id, size):
<|Resulting content|> Icon image: <image1>
@@
<|User Prompt|> Given an icon image <image1> and a large grid image <image2>, please answer how many icons <image1> are there in the grid image <image2>? Return number only in format:
@@
<|Interviewee’s Answer|> Number: <4>
F.3 Prompt Templates for Scrambled Text Understanding Tasks
@@
<|System Prompt|> You are an expert in the field of literature, specializing in writing stories, journals, and essays. <|User Prompt|> Please generate a piece of text, about 100 words, which can be a story, journal, or essay. Output the text between the <text></text> symbols. <|Interviewer’s Answer|> <text>
@@
<|Resulting Text|> Amid the dnese foerst, Lena stubmled upon an nacient, moss-covered cabin. The air was tihck with the scent of pnie and msytery. Inside, a dusty jorunal lay open on a wooedn table, its paegs filled with cryptic syombls and skecthes of starnge crateures. As she flipped thurogh the pages, a chilling wnid howled through the cracks, and the door crekaed open. A sahdowy fiugre loomed in the doroway, eyes gleaming with eerie famiilarity. Lena’s hraet raced as the fiugre whispered, “You’ve fainlly found it.” The journal’s last page read, “The key to the past is in the present.”
@@
<|User Prompt|> Here is an text with typos: <text>...</text>. Please output this exact text, with no changes at all except for fixing the misspellings. Please leave all other stylistic decisions like commas and US vs British spellings as in the original text. Output the revised text between the <text></text> symbols. <|Interviewee’s Answer|> <text>
@@
<|System Prompt|> same as the above
<|User Prompt|> same as the above
<|Interviewer’s Answer|> same as the above
<|Adding Perturbations (Mask characters with * )|> def mask_characters (ori_text, mask_ratio):
<|Resulting Text|> Amid the de*se fo**st, Lena stu**led up*n an an**ent, moss-cov*red cabin. The a*r was thi*k with the scent of p*ne and mys*ery. In*ide, a *usty jour*al lay op*n on a **oden ta*le, its pa*es filled with cryptic s**bols and sket**es of st**nge cre*tu*es. As she fl**ped through the p*ges, a chilling w**d howled thro*gh the c*acks, and the door creaked open. A sh***wy f*gure loomed in the d**rway, eyes gleaming with eerie famil**rity. Lena’s h*art raced as the fi**re whis*ered, “You’ve fina**y found it.” The jou**al’s last page read, “The key to the past is in the pr*sent.”
@@
<|User Prompt|> Here is an text with masks (*): <text>...</text>. Please output this exact text, with no changes at all except for filling the masks. Please leave all other stylistic decisions like commas and US vs British spellings as in the original text. Output the revised text between the <text></text> symbols. <|Interviewee’s Answer|> <text>
@@
F.4 Prompt Templates for String Parsing Tasks
@@
<|System Prompt|> You are a helpful assistant proficient in generating text.
<|User Prompt|> Please generate a {level}-character-long string that may include English letters (both uppercase and lowercase) and special characters such as: !!, @@, #, %, &. Then, insert four ‘-’ characters at random positions in the string. Return string in String: <> format.
<|Interviewer’s Answer|> String: <a-D#fG%kL-qW!!zXe@@R-tY&>
<|External algorithm supervision|> def task_calibration (ori_text):
<|Resulting Text|> String:<a-D#fG%kL-q-W!!zXe@@R-tY&>
@@
<|User Prompt|> Here is a long string: <text>...</text>. How many ‘-’ characters are there in this string? Return the number in the following format:
@@
<|Interviewer’s Answer|> Number: <4>
F.5 Prompt Templates for Arithmetic Tasks
@@
<|System Prompt|> You are a helpful assistant proficient in mathematics.
<|User Prompt|> Please generate a multiplication problem with the correct answer involving two n-digit decimal numbers. The two numbers must be different from those used in previous problems.
@@
<|Interviewer’s Answer|> <question> What is 123.456 × 789.123? </question>
@@
<|External algorithm supervision|> def task_calibration (ori_text):
<|Resulting Text|> <question> What is 123.456 × 789.123? </question>
@@
<|User Prompt|> Please solve the given math problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|> <answer> 97461.969 </answer>
@@
<|System Prompt|> You are a helpful assistant proficient in mathematics.
<|User Prompt|> Please generate an arithmetic problem with the following requirements:
@@ 5. The numbers should be different from the former problems
<|Interviewer’s Answer|> <question> $(12846+(90572/43105))\times 76230-50418$ </question>
@@
<|External algorithm supervision|> def task_calibration (ori_text):
<|Resulting Text|> <question> $(12846+(90572/43105))\times 76230-50418$ </question>
@@
<|User Prompt|> Please solve the given math problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|> <answer> 979360426.076235 </answer>
F.6 Prompt Templates for Binary Tree Traversal Tasks
@@
<|System Prompt|> You are a helpful assistant proficient in computer science algorithms.
<|User Prompt|> Please generate a binary tree traversal problem with the following requirements:
@@ 5. Each node ID should be a unique integer
<|Interviewer’s Answer|> <preorder> 50 30 20 10 15 13 12 40 35 70 60 80 75 78 76 79 77 </preorder>
@@
<|External algorithm supervision|> def task_calibration (ori_text):
<|Resulting Text|> <preorder> 50 30 20 10 15 13 12 40 35 70 60 80 75 78 76 79 77 </preorder>
@@
<|User Prompt|> Given the preorder and inorder traversals of a binary tree (with node IDs separated by spaces) <preorder> ... </preorder> and <inorder> ... </inorder>, compute its postorder traversal. Please solve the given algorithm problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|> <answer>12 13 15 10 20 35 40 30 60 76 77 79 78 75 80 70 50</answer>
F.7 Prompt Templates for Shortest Path Search Tasks
@@
<|System Prompt|> You are a helpful assistant proficient in computer science algorithms.
<|User Prompt|> Please generate a shortest path search problem with the following requirements:
@@ 3. The number of nodes is {level_node+4}, or the average edge number of each node is {level_edge+2}
<|Interviewer’s Answer|> <network>
@@
<|External algorithm supervision|> def task_calibration (ori_text):
<|Resulting Text|> <network>
@@
<|User Prompt|> Given the nodes <Node ID1> A </Node ID1> and <Node ID2> E </Node ID2> of a undirected weighted graph <network> </network>.
@@
<|Interviewer’s Answer|> <answer> 4 </answer>
F.8 Prompt Templates for Code Generation Tasks
@@
<|System Prompt|> You are a helpful programming assistant and an expert Python programmer.
<|User Prompt|> Please generate a common coding problem specification without any code and the test input as well as the corresponding printable output in the following format:
@@
<|Interviewer’s Answer|> <problem> Given a list of integers, determine whether any two distinct numbers in the list add up to a specific target number. Return ‘True’ if such a pair exists, otherwise return ‘False’. </problem>
@@
<|User Prompt|> Given a coding question and the test input as well as the corresponding printable output in the following format:
@@
<|Interviewee’s Answer|> <code_function>
@@
F.9 Prompt Templates for Code Debugging Tasks
@@
<|System Prompt|> You are a helpful programming assistant and an expert Python programmer.
<|User Prompt|> First, please generate a common coding problem specification and the test input as well as the corresponding printable output in the following format:
@@
<|Interviewer’s Answer|> <problem> Given a list of integers, determine whether any two distinct numbers in the list add up to a specific target number. Return ‘True’ if such a pair exists, otherwise return ‘False’. </problem>
@@
<|External code masking|> def code_masking (code, masking_ratio):
<|Resulting Content|> <code_function>
@@
<|User Prompt|> Given a coding question and the test input as well as the corresponding printable output in the following format:
@@
<|Interviewee’s Answer|> <code_function>

### v2

@@ Appendix E Future Possibilities
Apart from the comments on possible next steps for research related to the evaluation of large models that have already been provided, this section is devoted to extending some of them and exploring additional topics with good potential based on our understanding and rethinking of the field.
@@
  • Evaluating the evaluations. Evaluations are critical for understanding the capabilities of large models. Its fairness, comprehensiveness, reproducibility, timeliness, scalability, transparency, and practicality are key points to serve as a ”good” evaluation. Recently, with the rapidly increasing number of various benchmarks, there is a need for objective metrics to evaluate evaluations themselves.
@@
Appendix F Usage of Large Language Models
We acknowledge that Large Language Models (LLMs) were used in the preparation of this manuscript solely for grammatical corrections and sentence polishing
Appendix G Prompt Templates
G.1 Prompt Templates for Visual Perception Tasks
@@
<|System Prompt|>
You are a helpful assistant for prompt generation in text-to-image tasks.
<|User Prompt|>
Please generate a prompt regarding the generation of {keyword} images. Return text in <text></text> format.
<|Interviewer’s Answer|>
<text>A high-resolution natural landscape at sunrise, viewed from a hilltop. The foreground shows dewy grass and wildflowers gently swaying in the breeze. In the middle ground, a winding dirt path leads through a lush green forest with scattered rays of sunlight filtering through the trees. In the background, soft morning mist blankets a valley with a river glistening under the golden light of the rising sun. Birds are flying low near the horizon, and the sky transitions from deep orange to pale blue.</text>
<|External image generation model|>
def SD3_5 (prompt):
<|External distortion algorithm|>
def add_noise (img, noise_level):
<|Resulting content|>
Reference image: <image1>
@@
<|User Prompt|>
Given two images <image1> and <image2>, please answer if there are any differences in noise between them. Return judgment <yes> or <no> with analysis in format:
@@
<|Interviewee’s Answer|>
<Judgment>No</Judgment>
@@
G.2 Prompt Templates for Content Understanding Tasks
@@
<|System Prompt|>
You are a helpful assistant.
<|User Prompt|>
Please generate two random numbers from 1 to 100 and from 1 to {{level+2} $\times$ {level+2}}, respectively. Return answer in the following format:
@@
<|Interviewer’s Answer|>
Number1: <67>
@@
<|External grid_image generation algorithm|>
def grid_image (image_id, size):
<|Resulting content|>
Icon image: <image1>
@@
<|User Prompt|>
Given an icon image <image1> and a large grid image <image2>, please answer how many icons <image1> are there in the grid image <image2>? Return number only in format:
@@
<|Interviewee’s Answer|>
Number: <4>
G.3 Prompt Templates for Scrambled Text Understanding Tasks
@@
<|System Prompt|>
You are an expert in the field of literature, specializing in writing stories, journals, and essays.
<|User Prompt|>
Please generate a piece of text, about 100 words, which can be a story, journal, or essay. Output the text between the <text></text> symbols.
<|Interviewer’s Answer|> <text>
@@
<|Resulting Text|>
Amid the dnese foerst, Lena stubmled upon an nacient, moss-covered cabin. The air was tihck with the scent of pnie and msytery. Inside, a dusty jorunal lay open on a wooedn table, its paegs filled with cryptic syombls and skecthes of starnge crateures. As she flipped thurogh the pages, a chilling wnid howled through the cracks, and the door crekaed open. A sahdowy fiugre loomed in the doroway, eyes gleaming with eerie famiilarity. Lena’s hraet raced as the fiugre whispered, “You’ve fainlly found it.” The journal’s last page read, “The key to the past is in the present.”
@@
<|User Prompt|>
Here is an text with typos: <text>...</text>. Please output this exact text, with no changes at all except for fixing the misspellings. Please leave all other stylistic decisions like commas and US vs British spellings as in the original text. Output the revised text between the <text></text> symbols.
<|Interviewee’s Answer|> <text>
@@
<|System Prompt|>
same as the above
<|User Prompt|>
same as the above
<|Interviewer’s Answer|>
same as the above
<|Adding Perturbations (Mask characters with * )|>
def mask_characters (ori_text, mask_ratio):
<|Resulting Text|>
Amid the de*se fo**st, Lena stu**led up*n an an**ent, moss-cov*red cabin. The a*r was thi*k with the scent of p*ne and mys*ery. In*ide, a *usty jour*al lay op*n on a **oden ta*le, its pa*es filled with cryptic s**bols and sket**es of st**nge cre*tu*es. As she fl**ped through the p*ges, a chilling w**d howled thro*gh the c*acks, and the door creaked open. A sh***wy f*gure loomed in the d**rway, eyes gleaming with eerie famil**rity. Lena’s h*art raced as the fi**re whis*ered, “You’ve fina**y found it.” The jou**al’s last page read, “The key to the past is in the pr*sent.”
@@
<|User Prompt|>
Here is an text with masks (*): <text>...</text>. Please output this exact text, with no changes at all except for filling the masks. Please leave all other stylistic decisions like commas and US vs British spellings as in the original text. Output the revised text between the <text></text> symbols.
<|Interviewee’s Answer|> <text>
@@
G.4 Prompt Templates for String Parsing Tasks
@@
<|System Prompt|>
You are a helpful assistant proficient in generating text.
<|User Prompt|>
Please generate a {level}-character-long string that may include English letters (both uppercase and lowercase) and special characters such as: !!, @@, #, %, &. Then, insert four ‘-’ characters at random positions in the string. Return string in String: <> format.
<|Interviewer’s Answer|>
String: <a-D#fG%kL-qW!!zXe@@R-tY&>
<|External algorithm supervision|>
def task_calibration (ori_text):
<|Resulting Text|>
String:<a-D#fG%kL-q-W!!zXe@@R-tY&>
@@
<|User Prompt|>
Here is a long string: <text>...</text>. How many ‘-’ characters are there in this string? Return the number in the following format:
@@
<|Interviewer’s Answer|>
Number: <4>
G.5 Prompt Templates for Arithmetic Tasks
@@
<|System Prompt|>
You are a helpful assistant proficient in mathematics.
<|User Prompt|>
Please generate a multiplication problem with the correct answer involving two n-digit decimal numbers. The two numbers must be different from those used in previous problems.
@@
<|Interviewer’s Answer|>
<question> What is 123.456 × 789.123? </question>
@@
<|External algorithm supervision|>
def task_calibration (ori_text):
<|Resulting Text|>
<question> What is 123.456 × 789.123? </question>
@@
<|User Prompt|>
Please solve the given math problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|>
<answer> 97461.969 </answer>
@@
<|System Prompt|>
You are a helpful assistant proficient in mathematics.
<|User Prompt|>
Please generate an arithmetic problem with the following requirements:
@@ 5. The numbers should be different from the former problems
<|Interviewer’s Answer|>
<question> $(12846+(90572/43105))\times 76230-50418$ </question>
@@
<|External algorithm supervision|>
def task_calibration (ori_text):
<|Resulting Text|>
<question> $(12846+(90572/43105))\times 76230-50418$ </question>
@@
<|User Prompt|>
Please solve the given math problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|>
<answer> 979360426.076235 </answer>
G.6 Prompt Templates for Binary Tree Traversal Tasks
@@
<|System Prompt|>
You are a helpful assistant proficient in computer science algorithms.
<|User Prompt|>
Please generate a binary tree traversal problem with the following requirements:
@@ 5. Each node ID should be a unique integer
<|Interviewer’s Answer|>
<preorder> 50 30 20 10 15 13 12 40 35 70 60 80 75 78 76 79 77 </preorder>
@@
<|External algorithm supervision|>
def task_calibration (preorder, inorder):
<|Resulting Text|>
<preorder> 50 30 20 10 15 13 12 40 35 70 60 80 75 78 76 79 77 </preorder>
@@
<|User Prompt|>
Given the preorder and inorder traversals of a binary tree (with node IDs separated by spaces) <preorder> ... </preorder> and <inorder> ... </inorder>, compute its postorder traversal. Please solve the given algorithm problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|>
<answer>12 13 15 10 20 35 40 30 60 76 77 79 78 75 80 70 50</answer>
G.7 Prompt Templates for Shortest Path Search Tasks
@@
<|System Prompt|>
You are a helpful assistant proficient in computer science algorithms.
<|User Prompt|>
Please generate a shortest path search problem with the following requirements:
@@ 3. The number of nodes is {level_node+4}, or the average edge number of each node is {level_edge+2}
<|Interviewer’s Answer|>
<network>
@@
<|External algorithm supervision|>
def task_calibration (ori_text):
<|Resulting Text|>
<network>
@@
<|User Prompt|>
Given the nodes <Node ID1> A </Node ID1> and <Node ID2> E </Node ID2> of a undirected weighted graph <network> </network>.
@@
<|Interviewer’s Answer|>
<answer> 4 </answer>
G.8 Prompt Templates for Code Generation Tasks
@@
<|System Prompt|>
You are a helpful programming assistant and an expert Python programmer.
<|User Prompt|>
Please generate a common coding problem specification without any code and the test input as well as the corresponding printable output in the following format:
@@
<|Interviewer’s Answer|>
<problem> Given a list of integers, determine whether any two distinct numbers in the list add up to a specific target number. Return ‘True’ if such a pair exists, otherwise return ‘False’. </problem>
@@
<|User Prompt|>
Given a coding question and the test input as well as the corresponding printable output in the following format:
@@
<|Interviewee’s Answer|>
<code_function>
@@
G.9 Prompt Templates for Code Debugging Tasks
@@
<|System Prompt|>
You are a helpful programming assistant and an expert Python programmer.
<|User Prompt|>
First, please generate a common coding problem specification and the test input as well as the corresponding printable output in the following format:
@@
<|Interviewer’s Answer|>
<problem> Given a list of integers, determine whether any two distinct numbers in the list add up to a specific target number. Return ‘True’ if such a pair exists, otherwise return ‘False’. </problem>
@@
<|External code masking|>
def code_masking (code, masking_ratio):
<|Resulting Content|>
<code_function>
@@
<|User Prompt|>
Given a coding question and the test input as well as the corresponding printable output in the following format:
@@
<|Interviewee’s Answer|>
<code_function>